OpenStudy (anonymous):

Simplify the following expression: 21(x-4)+2

OpenStudy (anonymous):

??????????????????????????????????

OpenStudy (wendy.ivette11714):

\[21(x-4)+2\] Distribute:\[=(21)(x)+(21)(−4)+2\] \[=21x+−84+2\] Combine like terms: \[=21x+−84+2\] \[=(21x)+(−84+2)\] \[=21x+−82\]

OpenStudy (anonymous):

So you just add 21+-82?

OpenStudy (anonymous):

What does the x mean?

OpenStudy (anonymous):

oh nvm....

OpenStudy (wendy.ivette11714):

21x + -82 is the answer

OpenStudy (anonymous):

yeah I see that now... thanks :)

OpenStudy (wendy.ivette11714):

yw :)

OpenStudy (anonymous):

well for my answer it shows - not +...